# An Atomic Cluster Expansion Potential for Twisted Multilayer Graphene

### v1

@@ Abstract
Twisted multilayer graphene, characterized by its moiré patterns arising from inter-layer rotational misalignment, serves as a rich platform for exploring quantum phenomena. While first-principles calculations are computationally prohibitive and empirical interatomic potentials often lack accuracy, machine-learning interatomic potentials (MLIPs) present a promising alternative, offering (near-)DFT accuracy at a significantly reduced computational cost. Despite their success in two-dimensional monolayer materials, MLIPs remain under-explored in twisted multilayer graphene systems. In this work, we develop an Atomic Cluster Expansion (ACE) potential for simulating twisted multilayer graphene and test it on a range of simulation tasks. We propose an approach to construct training and test datasets that incorporate all possible twist angles and local stacking, including incommensurate ones. To achieve this, we generate configurations with periodic boundary conditions suitable for DFT calculations, and then introduce an internal twist and shift within those supercell structures. We further refine the dataset through active learning filtering, guided by Bayesian uncertainty quantification. Our model is validated for accuracy and robustness through a wide range of numerical tests.
@@ I Introduction
Multilayer graphene has garnered significant attention due to its unique electronic properties and potential applications in next-generation devices [1, 2]. When graphene layers are stacked with a slight rotational misalignment, known as a twist angle, they form moiré patterns, leading to novel physical phenomena [3, 4, 5, 6]. Understanding these phenomena is crucial for advancing condensed matter physics and developing materials with tunable electronic characteristics. The ability to manipulate these properties through the twist angle makes the twisted graphene systems a versatile platform for exploring quantum mechanical effects. Consequently, research into twisted bilayer and multilayer graphene holds great promise for innovations in nanoelectronics, photonics, and quantum computing.
@@
The Atomic Cluster Expansion (ACE), introduced in [16], is a particularly flexible and theoretically sound MLIP approach. ACE offers a balance between cost and accuracy and has been successfully applied across a wide range of tasks, particularly in materials simulation [21, 22, 23, 24, 25, 26, 27, 28]. Linear variants of the ACE model have proven to be both data and computationally efficient, making them particularly useful in active learning (AL) workflows [25]. Linearity also facilitates sensitivity analysis and paves the way for reliable uncertainty quantification. Throughout this work, we adopt the ACE method, but the methods and benchmarks we introduce are likely applicable to the vast majority of MLIPs frameworks.
@@
Although MLIPs have been extensively utilized in various fields, their application to twisted multilayer graphene systems remains largely unexplored. Previous studies have developed MLIPs for other two-dimensional (2D) materials, such as MoS2, monolayer graphene, buckled silicone, and quaternary transition metal dichalcogenide (TMD) alloy, to investigate thermal conductivity and phononic properties [29, 30, 31, 32, 33]. For instance, Magorrian et al. [34] focused on atomistic relaxation in twisted bilayers of InSe, while Jung et al. [35] explored mechanical and fracture dynamics in monolayer graphene. Yin et al. [36] applied machine-learning force fields for membrane design, and Wilson et al. [37] used batch active learning to construct MLIPs, emphasizing phonon dispersion and phase transition temperature calculations for monolayer GeSe. More recently, Bao et al. [38] leveraged transfer learning to predict the electronic structure of twisted MoTe2, while a closely related study [39] developed a framework for constructing MLIPs for moiré systems, specifically MX2 (M = Mo, W; X = S, Se, Te) materials. Georgaras et al. [40] further advanced this direction by proposing a split MLIP architecture and a curated dataset strategy that explicitly distinguishes intralayer and interlayer interactions, leading to improved accuracy in modeling moiré materials. However, these studies primarily focus on relatively simple physical properties such as relaxed lattice constants, phonon spectra, and band structures. To the best of our knowledge, no MLIP has been specifically developed for twisted multilayer graphene systems, nor has there been a comprehensive benchmark of such systems, which serves as the primary motivation for our work.
@@
In this study, we develop a training and benchmark protocol for twisted multilayer graphene systems and apply it to fitting an ACE model. A key challenge in developing an MLIP is the selection of small-scale training structures. For twisted multilayer graphene this is particularly challenging since most twist angles of interest cannot be represented in a small periodic structure accessible to DFT methods. A novel contribution of our work to overcome this is to employ locally twisted training structures. This provides a systematic method to generate training configurations with periodic boundary conditions that are suitable for DFT calculations across all twist angles. We combine this approach with Bayesian active learning and filtering techniques that sub-select configurations based on uncertainties derived from Bayesian linear regression.
@@ II.1 Multilayer twisted graphene
Each unit cell of graphene has two carbon atoms with shifts $b_{A}=(0,0)^{T}$ and $b_{B}=a_{0}(0,\sqrt{3}/3)^{T}$, labeled “A site” and “B site” in Figure 1. For our numerical results we used the lattice constant $a_{0}=2.47$. Twisted bilayer graphene is formed by stacking two layers of graphene on top of each other with a relative twist. The two layers’ fundamental matrices are then:
@@
The misalignment between the two layers caused by the twist angle produces a large-scale moiré pattern. The resulting bilayer structure is generally incommensurate and only exhibits approximate periodicity. Exact periodicity is achieved only at specific twist angles where $\cos\theta$ satisfies a Diophantine condition [41, 42]. The moiré unit cell has the form [43, 44, 45]:
@@
The moiré superlattice length depends on the rotation angle between two layers $\theta$, growing as $\sim 1/\theta$ as $\theta$ decreases. Techniques that rely on periodic boundary conditions or supercell approximations require impractically large domains for numerical purposes. This effect is even more severe in the multi-layer case [46, 47].
The large length scale of moiré patterns and lack of periodicity make numerical simulation of many interesting phenomena in the small-angle regime challenging. A wide variety of strongly correlated electronic phases of matter such as superconductivity [48] and fractional quantum anomalous Hall (fQAHE) states [49, 50] can emerge. Mechanical properties, which are the primary focus of the present paper, including structural relaxation [44, 47, 51, 5, 52, 45, 53, 54], phonons [55, 56, 57], and defects [58, 59], play a critical role in understanding the electronic behavior of these moiré systems. For example, mechanical relaxation can form large triangular or hexagonal domains [60], modifying electronic structures and shifting the twist angle at which fQAHE is expected to appear. Additionally, strong electron-phonon coupling, which has been measured [61], may contribute to the observed moiré superconductivity. Defects are also prevalent in these moiré systems and their impact on electronic properties is not yet understood.
Mechanical relaxation of multilayer graphene systems has in the past been performed using linear elasticity approximations or the REBO interatomic potential for the intralayer component and the generalized stacking fault energy or the Kolmogorov-Crespi interatomic potential [10] for the interlayer component [62, 44, 45, 47, 51, 52]. Although the Kolmogorov-Crespi interatomic potential [10] is a major improvement over two-body interatomic potentials such as Lennard-Jones, there are significant errors in the prediction of forces of some important local configurations (disregistries) when compared to DFT computations  [63]. An important perspective of our present work is to develop a methodology (in the simplified multi-layer graphene setting) that can be readily extended to systematically construct accurate interatomic potentials for more complex 2D materials such as the transition metal dichalcogenides (TMDs).
@@ II.2 The linear ACE framework
where $N_{\mathbf{A}}$ is the number of atoms in structure $\mathbf{A}$, $F(\mathbf{A};{\bf c})$ is a vector of forces, $\Sigma(\mathbf{A};\mathbf{c})$ is the virial stress of the ACE model, $\Gamma$ specifies the form of the regularizer (smoothness prior), and $\lambda$ is a scaling parameter that determines the relative weight of regularization. The weights $w_{E}$, $w_{F}$ and $w_{\Sigma}$ adjust the contributions of energy, forces and virial stresses, respectively. Within the Bayesian interpretation of ridge regression (also used as the solver, see the discussion below), $\Gamma$ defines a prior distribution for the model parameters $\mathbf{c}\sim\mathcal{N}({\bf 0},\Gamma^{-2})$. For details, see Appendix B and [64]. Throughout the present work, we choose an algebraic smoothness prior of strength $p=5$, as defined in [64, Eq.10].
@@ II.3.1 Locally twisted multi-layers
The procedure described in the foregoing paragraph can be extended to three or more alternating supercell layers. The main restriction remains the periodicity of every layer on the chosen supercell limiting the choices of global twist angles between layers. Here, we employ alternating twist angles where all adjacent layers are twisted relative to each other by alternating $\theta_{\text{sc}}$ or $-\theta_{\text{sc}}$, where $\theta_{\text{sc}}$ is the commensurate twist angle corresponding to the supercell. Let the origin again be the center of the commensurate supercell. In this case of alternating twist, we apply a similar deformation map
@@ II.3.2 Filtering the candidate dataset
The trilayer, quadrilayer and pentalayer systems were also all generated using the local twisting method. The commensurate twist angles used were the same as in the bilayer case (no twist, $4.41^{\circ}$, $16.4^{\circ}$). Since any comprehensive sampling of the alternating trilayer, quadrilayer and pentalayer would yield too large of a dataset, we randomly selected disregistries, interlayer distances and local twist angle differences. We obtained 500 unrattled and 500 rattled (randomly perturbed) samples for the trilayer, quadrilayer and pentalayer datasets. The disregistries were randomly selected uniformly over the unit cell. The interlayer distances were uniformly sampled between 2.7 Å and 4 Å independently for each pair of layers. The local twist angle differences ($\theta_{j}$’s from the local twisting method) were uniformly chosen between $-15^{\circ}$ and $15^{\circ}$. For the rattled samples, the rattling standard of deviation was selected uniformly from 00 to $0.15$ times the AB distance. We could have additionally exploited the symmetries when all layers are rotated or shifted together, however, the hyperactive learning process will bias sampling away from these symmetry directions regardless.
@@
    Choose ACE hyperparameters: Correlation order $\nu_{\rm max}=4$ (body-order five); Polynomial total degree 20; Cutoff radius $r_{\rm cut}=5.5$; regression weights $w_{E}/w_{F}/w_{V}=10.0/1.0/1.0$; Committee size $K=20$. The resulting model has approximately 5k parameters.
@@
This filtering strategy effectively reduced the dataset to approximately 5% of the original candidates while achieving strong generalization between training and testing errors, as we demonstrate in the next section.
@@ II.3.3 Fine-tuning
It is all but impossible to fit an MLIP that has good quantitative accuracy across all possible simulation tasks for a given material. The models we report will be shown to have excellent accuracy in-distribution (near the training set) and behave robustly with good qualitative accuracy out-of-distribution. In order to obtain quantitatively accurate result on new simulation tasks, it may be necessary to extend the training data in a way that is representative of the new tasks, or within an active learning procedure. Given the low cost of fitting a linear ACE potential it is straightforward to then simply refit the model to the expanded dataset, normally with the same hyperparameters. We will demonstrate this on a limited set of out-of-distribution tests. We call this procedure “fine-tuning” to indicate the analogy with fine-tuning foundation models [66].
@@ III.1 Fitting Accuracy
Figure 4 provides a more detailed description of the force accuracy of the ACE potentials. In each column, the three subplots represent the predicted forces versus the DFT reference forces for monolayer, twisted bilayer, and twisted trilayer configurations, respectively. The results demonstrate that training only on mono-layer configurations (ACE (mo)) does not yield any acceptable generalization to bi- and tri-layers. This is unsurprising since the cutoff radius for the ACE potentials is larger than the interlayer distance. Training on only bilayer configurations results in moderately accurate generalization to both monolayer and twisted trilayer configurations. Note in particular that the generic pacemaker model [28] has far better accuracy on all multi-layers except bilayers.
@@ III.2.2 Defect formation energies
The previous section demonstrates the stability of geometry optimization. Next, we test quantitative accuracy of the predicted minimizer by comparing predicted defect formation energies against their corresponding DFT values. These tests assess out-of-distribution performance, since the training set does not contain defective structures. The defect-free system is a $4.41^{\circ}$ degree twisted bilayer structure containing 676 atoms. Interatomic forces are relaxed in all the systems using the trained interatomic potentials. The reference DFT energetics are computed on the relaxed structures using the trained potentials. Further relaxation of DFT interatomic forces is not conducted. Given that the model ACE (gt) trained only on globally twisted configurations does not perform well, we limit these tests to the ACE (lt), ACE (gl) and ACE ($\ast$) models. The results are summarized in Table 5. We observe that the ACE ($\ast$) model (generated using the active learning strategy) provides significantly higher accuracy on all tests, with the exception of a single test where the pacemaker potential achieves slightly better accuracy. The models trained solely on locally twisted configurations and those trained on combined data all show relatively crude accuracy, comparable with the general purpose pacemaker potential.
Table 5: Point defect formation energies in bilayer graphene with a twist angle of $4.41^{\circ}$ for three ACE potentials. The relative error is given in the bracket. Entries within ±$\pm 5\%$ of the reference values and the best-performing results across all methods are highlighted.

### v2

@@ Abstract
Twisted multilayer graphene, characterized by its moiré patterns arising from inter-layer rotational misalignment, serves as a rich platform for exploring quantum phenomena. Machine learning interatomic potentials (MLIPs) are a promising approach to model such systems. Our work develops a method to generate training and test datasets for fitting MLIPs that capture all possible misalignments but remain small-scale to facilitate efficient data generation and parameter estimation. To achieve this, we generate configurations with periodic boundary conditions suitable for DFT calculations, and then introduce an internal twist and shift within those supercell structures. Using this technique, supplemented with an active learning workflow, we fit an Atomic Cluster Expansion potential for simulating twisted multilayer graphene and test it for accuracy and robustness on a range of simulation tasks.
@@ I Introduction
Multilayer graphene has garnered significant attention due to its unique electronic properties and potential applications in next-generation devices [1, 2]. When graphene layers are stacked with a slight rotational misalignment, known as a twist angle, they form moiré patterns, leading to novel physical phenomena [3, 4, 5, 6, 7]. Understanding these phenomena is crucial for advancing condensed matter physics and developing materials with tunable electronic characteristics. The ability to manipulate these properties through the twist angle makes the twisted graphene systems a versatile platform for exploring quantum mechanical effects. Consequently, research into twisted bilayer and multilayer graphene holds great promise for innovations in nanoelectronics, photonics, and quantum computing.
@@
The Atomic Cluster Expansion (ACE), introduced in [17], is a particularly flexible and theoretically sound MLIP approach. ACE offers a balance between cost and accuracy and has been successfully applied across a wide range of tasks, particularly in materials simulation [22, 23, 24, 25, 26, 27, 28]. Linear variants of the ACE model have proven to be both data and computationally efficient, making them particularly useful in active learning (AL) workflows [25]. Linearity also facilitates sensitivity analysis and paves the way for reliable uncertainty quantification. Throughout this work, we adopt the ACE method, but the methods and benchmarks we introduce are likely applicable to the vast majority of MLIPs frameworks.
@@
Although MLIPs have been extensively utilized in various fields, their application to twisted multilayer graphene systems remains largely unexplored. Previous studies have developed MLIPs for other two-dimensional (2D) materials, such as MoS2, monolayer graphene, buckled silicone, and quaternary transition metal dichalcogenide (TMD) alloy, to investigate thermal conductivity and phononic properties [29, 30, 31, 32, 33]. For instance, Magorrian et al. [34] focused on atomistic relaxation in twisted bilayers of InSe, while Jung et al. [35] explored mechanical and fracture dynamics in monolayer graphene. Yin et al. [36] applied machine-learning force fields for membrane design, and Wilson et al. [37] used batch active learning to construct MLIPs, emphasizing phonon dispersion and phase transition temperature calculations for monolayer GeSe. More recently, Bao et al. [38] leveraged transfer learning to predict the electronic structure of twisted MoTe2, while a closely related study [39] developed a framework for constructing MLIPs for moiré systems, specifically MX2 (M = Mo, W; X = S, Se, Te) materials. Georgaras et al. [40] further advanced this direction by proposing a split MLIP architecture and a curated dataset strategy that explicitly distinguishes intralayer and interlayer interactions, leading to improved accuracy in modeling moiré materials. These studies provide a general, theoretically and computationally attractive approach to predicting physical properties on homogeneous structures, such as relaxed lattice constants, phonon spectra, and band structures.
@@
In the present study, we develop methods to expand the training set and extensive benchmarking of MLIPs for twisted multilayer systems in- and out-of-distribution; to be used in a stand-alone fashion or integrated into more general frameworks such as DPmoire [39]. Specifically, we develop a training and benchmark protocol for twisted multilayer graphene systems and apply it to fitting an ACE model. A key challenge in developing an MLIP is the selection of small-scale training structures. For twisted multilayer graphene this is particularly challenging since most twist angles of interest cannot be represented in a small periodic structure accessible to DFT methods. A novel contribution of our work to overcome this is to employ locally twisted training structures. This provides a systematic method to generate training configurations with periodic boundary conditions that are suitable for DFT calculations across all twist angles. We combine this approach with Bayesian active learning and filtering techniques that sub-select configurations based on uncertainties derived from Bayesian linear regression.
@@ II.1 Multilayer twisted graphene
Each unit cell of graphene has two carbon atoms with shifts $b_{A}=(0,0)^{T}$ and $b_{B}=a_{0}(0,\sqrt{3}/3)^{T}$, labeled “A site” and “B site” in Figure 1. For our numerical results we used the lattice constant $a_{0}=2.47$ Å, which is taken from experimental measurements as reported in [41]. Twisted bilayer graphene is formed by stacking two layers of graphene on top of each other with a relative twist. The two layers’ fundamental matrices are then:
@@
The misalignment between the two layers caused by the twist angle produces a large-scale moiré pattern. The resulting bilayer structure is generally incommensurate and only exhibits approximate periodicity. Exact periodicity is achieved only at specific twist angles where $\cos\theta$ satisfies a Diophantine condition [41, 42]. The moiré unit cell has the form [43, 44]:
@@
The moiré superlattice length depends on the rotation angle between two layers $\theta$, growing as $\sim 1/\theta$ as $\theta$ decreases. Techniques that rely on periodic boundary conditions or supercell approximations require impractically large domains for numerical purposes. This effect is even more severe in the multi-layer case [45].
The large length scale of moiré patterns and lack of periodicity make numerical simulation of many interesting phenomena in the small-angle regime challenging. A wide variety of strongly correlated electronic phases of matter such as superconductivity [46] and fractional quantum anomalous Hall (fQAHE) states [47, 48] can emerge. Mechanical properties, which are the primary focus of the present paper, including structural relaxation [49, 50, 51, 52, 44, 53], phonons [54, 55], and defects, play a critical role in understanding the electronic behavior of these moiré systems. For example, mechanical relaxation can form large triangular or hexagonal domains [49], modifying electronic structures and shifting the twist angle at which fQAHE is expected to appear. Additionally, strong electron-phonon coupling, which has been measured [56], may contribute to the observed moiré superconductivity. Defects are also prevalent in these moiré systems and their impact on electronic properties is not yet understood.
Mechanical relaxation of multilayer graphene systems has in the past been performed using linear elasticity approximations or the REBO interatomic potential for the intralayer component and the generalized stacking fault energy or the Kolmogorov-Crespi interatomic potential [11] for the interlayer component [57, 44, 45, 52]. Although the Kolmogorov-Crespi interatomic potential [11] is a major improvement over two-body interatomic potentials such as Lennard-Jones, there are significant errors in the prediction of forces of some important local configurations (disregistries) when compared to DFT computations  [58]. An important perspective of our present work is to develop a methodology (in the simplified multi-layer graphene setting) that can be readily extended to systematically construct accurate interatomic potentials for more complex 2D materials such as the transition metal dichalcogenides (TMDs).
@@ II.2 The linear ACE framework
where $N_{\mathbf{A}}$ is the number of atoms in structure $\mathbf{A}$, $F(\mathbf{A};{\bf c})$ is a vector of forces, $\Sigma(\mathbf{A};\mathbf{c})$ is the virial stress of the ACE model, $\Gamma$ specifies the form of the regularizer (smoothness prior), and $\lambda$ is a scaling parameter that determines the relative weight of regularization. The weights $w_{E}$, $w_{F}$ and $w_{\Sigma}$ adjust the contributions of energy, forces and virial stresses, respectively. Within the Bayesian interpretation of ridge regression (also used as the solver, see the discussion below), $\Gamma$ defines a prior distribution for the model parameters $\mathbf{c}\sim\mathcal{N}({\bf 0},\Gamma^{-2})$. For details, see Appendix B and [59]. Throughout the present work, we choose an algebraic smoothness prior of strength $p=5$, as defined in [59, Eq.10].
@@ II.3.1 Locally twisted multi-layers
The procedure described in the foregoing paragraph can be extended to three or more alternating supercell layers. The main restriction remains the periodicity of every layer on the chosen supercell limiting the choices of global twist angles between layers. Here, we employ alternating twist angles where all adjacent layers are twisted relative to each other by alternating $\theta_{\text{sc}}$ or $\theta_{\text{sc}}$, where $\theta_{\text{sc}}$ is the commensurate twist angle corresponding to the supercell. Let the origin again be the center of the commensurate supercell. In this case of alternating twist, we apply a similar deformation map
@@ II.3.2 Filtering the candidate dataset
The trilayer, quadrilayer and pentalayer systems were also all generated using the local twisting method. The commensurate twist angles used were the same as in the bilayer case (no twist, $4.41^{\circ}$, $16.4^{\circ}$). Since any comprehensive sampling of the alternating trilayer, quadrilayer and pentalayer would yield too large of a dataset, we randomly selected disregistries, interlayer distances and local twist angle differences. We obtained 500 unrattled and 500 rattled (randomly perturbed) samples for the trilayer, quadrilayer and pentalayer datasets. The disregistries were randomly selected uniformly over the unit cell. The interlayer distances were uniformly sampled between 2.7 Å and 4 Å independently for each pair of layers. The local twist angle differences ($\theta_{j}$’s from the local twisting method) were uniformly chosen between $15^{\circ}$ and $15^{\circ}$. For the rattled samples, the rattling standard of deviation was selected uniformly from 0 to $0.15$ times the AB distance. We could have additionally exploited the symmetries when all layers are rotated or shifted together, however, the hyperactive learning process will bias sampling away from these symmetry directions regardless.
@@
    Choose ACE hyperparameters: Correlation order $\nu_{\rm max}=4$ (body-order five); Polynomial total degree 20; Cutoff radius $r_{\rm cut}=5.5$ Å; regression weights $w_{E}/w_{F}/w_{V}=10.0/1.0/1.0$; Committee size $K=20$. The resulting model has approximately 5k parameters.
@@
This filtering strategy effectively reduced the dataset to approximately 5% of the original candidates while achieving strong generalization between training and testing errors, as we demonstrate in the next section. The workflow is illustrated in Figure 4.
Figure 4: Workflow of the proposed framework, from dataset construction and active learning filtering to ACE model training and benchmark evaluations.
@@ II.3.3 Fine-tuning
It is all but impossible to fit an MLIP that has good quantitative accuracy across all possible simulation tasks for a given material. The models we report will be shown to have excellent accuracy in-distribution (near the training set) and behave robustly with good qualitative accuracy out-of-distribution. In order to obtain quantitatively accurate result on new simulation tasks, it may be necessary to extend the training data in a way that is representative of the new tasks, or within an active learning procedure. Given the low cost of fitting a linear ACE potential it is straightforward to then simply refit the model to the expanded dataset, normally with the same hyperparameters. We will demonstrate this on a limited set of out-of-distribution tests. We call this procedure “fine-tuning” to indicate the analogy with fine-tuning foundation models.
@@ III.1 Fitting Accuracy
Figure 5 provides a more detailed description of the force accuracy of the ACE potentials. In each column, the three subplots represent the predicted forces versus the DFT reference forces for monolayer, twisted bilayer, and twisted trilayer configurations, respectively. The results demonstrate that training only on mono-layer configurations (ACE (mo)) does not yield any acceptable generalization to bi- and tri-layers. This is unsurprising since the cutoff radius for the ACE potentials is larger than the interlayer distance. Training on only bilayer configurations results in moderately accurate generalization to both monolayer and twisted trilayer configurations. Note in particular that the generic pacemaker model [28] has far better accuracy on all multi-layers except bilayers. Our approach benefits from a training dataset specifically curated for twisted graphene systems such as local twist angle variations and stacking disorder. This task-specific dataset results in higher accuracy of the fitted linear ACE model on our specific tests. This accuracy improvement is likely independent of the chosen MLIP model.
@@ III.2.2 Defect formation energies
The previous section demonstrates the stability of geometry optimization. Next, we test quantitative accuracy of the predicted minimizer by comparing predicted defect formation energies against their corresponding DFT values. These tests assess out-of-distribution performance, since the training set does not contain defective structures. The defect-free system is a $4.41^{\circ}$ degree twisted bilayer structure containing 676 atoms. Interatomic forces are relaxed in all the systems using the trained interatomic potentials. The reference DFT energetics are computed on the relaxed structures using the trained potentials. Further relaxation of DFT interatomic forces is not conducted. Given that the model ACE (gt) trained only on globally twisted configurations does not perform well, we limit these tests to the ACE (lt), ACE (gl) and ACE ($\ast$) models. The results are summarized in Table 5. We observe that the ACE ($\ast$) model (generated using the active learning strategy) provides significantly higher accuracy on all tests, with the exception of a single test where the pacemaker potential achieves slightly better accuracy. The models trained solely on locally twisted configurations and those trained on combined data all show relatively crude accuracy, comparable with the general purpose pacemaker potential. To improve model transferability to defective configurations, we additionally included 50 near-equilibrium defect structures (vacancy, Stone–Wales, and interstitial types) in the training set. The resulting ACE (ft) shows consistently improved accuracy in predicting defect formation energies compared to models trained without such data.
Table 5: Point defect formation energies in bilayer graphene with a twist angle of $4.41^{\circ}$ for three ACE potentials. The relative error is given in the bracket. Entries within $\pm 5\%$ of the reference values and the best-performing results across all methods are highlighted.